OpenStudy (anonymous):

How do I solve the following equations for the value of x: 3x+=4x-10, 4x+10=6x-8, 3x+10=5x, 3x-11=8x-21, 3(x+8)=12

OpenStudy (welshfella):

look like there is a number missing form the first one but lets look at the second:- 4x + 10 = 6x - 8 we need to bring all the terms in x ( that is 4x and 6x) to one side of the '=' and the numbers on the other side

OpenStudy (anonymous):

the equation for the first one is not missing any numbers

OpenStudy (welshfella):

so first we subtract 4x from LHS (left hand side):- 4x - 4x + 10 = 6x - 4x - 9 4x - 4x = 0 and 6x - 4x = 2x so:- 10 = 2x - 9 now we need to get the 9 on LHS so since -9 + 9 = 0 we add 9 to both sides: 10 + 9 = 2x - 9 + 9 19 = 2x now if we divide both sides by 2 we get the value of x 19/2 = 2x / 2

OpenStudy (welshfella):

ok - with the first one I thought it might be because you have a + with nothing after it

OpenStudy (welshfella):

oops sorry - I made a typo in the second equation the last number is 8 not 9 so the last line will be 18/2 = 2x so you only have to divide 18 by 2 and you have found the value of x and the equation is solved

OpenStudy (welshfella):

do you follow the method ok?

OpenStudy (welshfella):

the first 4 equation can be solved in the same way eqaution 5 3(x + 8) = 12 you first have to distribute the 3 over the parentheses so its 3*x + 3*8 = 12 3x + 24 = 12

OpenStudy (anonymous):

Are you having difficulties solving this?

OpenStudy (anonymous):

can you help me solve the rest please I'm having little difficulties